Question 1:
A researcher is interested in knowing if there are differences between the incidence of accidents
among pilots of different experience levels. They were able to collect 62 responses from a random number of pilots. Among them, 17 new pilots were in accidents, 26 were not;; 13 experienced pilots were in accidents, 6 were not.

• Construct a contingency table.
• Run the Chi-­Square test for independence using StatCrunch.
• To access, click STAT>Tables>Contingency> With summary
• State your findings and their significance. Do not forget to give summary statistics in
APA format.

Question 2:
A researcher has found that pilots taking an over the counter antihistamine have poor reaction
times. They performed an experiment on 22 pilots: 11 took the medication and 11 did not. Their
reaction times to a command by Air Traffic Control was considered to be adequate or
inadequate. Among those taking the medication, 8 were rated inadequate and 3 adequate.
Among those not-­taking the medication, 10 were rated adequate and only 1 inadequate.

• Construct a contingency table.
• Run the Chi-­Square test for independence using StatCrunch.
• To access, click STAT>Tables>Contingency> With summary
• State your findings and their significance. Do not forget to give summary statistics in
APA format.
• What is the problem with this research study? Is the Chi-­Square test appropriate? Why
or why not? What could the researcher do to potentially rectify the problem
Answer #1

Solution:-

1)

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

H0: The null hypothesis states that the proportion of accidents by experienced drivers is identical to the proportion of accidents by non-experienced drivers.

Alternative hypothesis: At least one of the null hypothesis statements is false.

Formulate an analysis plan. For this analysis, the significance level is 0.05. Using sample data, we will conduct a chi-square test for independence.

Analyze sample data. Applying the chi-square test for independence to sample data, we compute the degrees of freedom, the expected frequency counts, and the chi-square test statistic. Based on the chi-square statistic and the degrees of freedom, we determine the P-value.

DF = (r - 1) * (c - 1) = (2 - 1) * (2 - 1)
D.F = 1
Er,c = (nr * nc) / n

Accidents Not Accidents Total New pilots Experiened Total 26 17 13 30 43 19 62 32

\chi ^{2}=\sum \left [\frac{(O_{r,c}-E_{r,c})^{2}}{E_{r,c}} \right ]

Χ2 = 4.40

where DF is the degrees of freedom.

The P-value is the probability that a chi-square statistic having 1 degrees of freedom is more extreme than 4.40.

We use the Chi-Square Distribution Calculator to find P(Χ2 > 4.40) = 0.036.

Interpret results. Since the P-value (0.036) is less than the significance level (0.05), we cannot accept the null hypothesis.